# Binary Stars Approaching Supermassive Black Holes: Hydrodynamics of Stellar Collisions, Mass Fallback and Partial TDEs

### v1

@@ Abstract
When binaries are injected into low-angular-momentum orbits around a central supermassive black hole (SMBH), various outcomes can occur, including binary tidal breakup, double stellar disruptions and stellar collision. We use hydrodynamical simulations to study stellar collisions triggered by binary-SMBH encounters, examining both head-on and grazing collisions in deep ($\beta_{b}=5$) and gentle ($\beta_{b}=0.6$) encounters, where $\beta_{b}$ is the ratio of the binary tidal disruption radius to the binary pericenter distance to the SMBH. Head-on collisions consistently result in appreciable mass loss ($\sim 5\%$) and a single merger remnant. Grazing collisions in deep encounters typically leave two strongly disturbed stars with comparable mass loss, while in gentle encounters, multiple collisions eventually produce a single remnant with minimal mass loss ($\lesssim 1\%$). All merger remnants feature extended envelopes, making them susceptible to partial tidal disruptions when they return to the SMBH. The morphology and orbital energy distribution of collision-induced debris differ significantly from those of tidal disruption event (TDE) debris of single stars. Approximately half of the collision-generated debris falls back onto the SMBH, exhibiting a distinct time evolution of the fallback rate. We suggest that such mass loss and fallback can generate electromagnetic flares that mimic weak TDEs.
@@
software: PHANTOM (Price et al., 2018), Sarracen (Harris & Tricco, 2023), Matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011).
@@ 1 Introduction
In Yu & Lai (2024) (hereafter Paper I), we used three-body scattering experiments to characterize the key dynamics of binary tidal breakups, double stellar disruptions, and stellar collisions induced by binary-SMBH encounters. We analyzed the collision probabilities of binaries with different orbital properties and concluded that the collision probability can reach a few to 10’s percent as the binary binary’s pericenter distance (relative to the SMBH) varies from outside the tidal sphere to within it. Additionally, we found that the impact velocities of the colliding stars are comparable to the escape velocity of the star.
@@
Our paper is organized as follows. In Section 2, we describe the setup of our numerical simulations, including the initial conditions and methodology. In Section 3, we present the results from our representative simulations. Finally, in Section 4, we summarize our findings and comment on the results of Paper I based on our hydrodynamical simulations.
@@ 2 Method and Setup
    $\alpha\equiv a_{b}/R_{\rm col}$, where $R_{\rm col}$ is the separation between the two stars at the moment of collision. For an equal-mass binary ($m_{1}=m_{2}$), $R_{\rm col}=2R_{\star}$.
@@ 2.1 Simulation Setup
Throughout this paper, we consider circular equal-mass binary, where each star modeled as a $\gamma=5/3$ polytrope, with mass $m_{\star}=1M_{\odot}$ and radius $R_{\star}=1R_{\odot}$. The binaries in our simulations are then created by shifting the centers of two stellar models; each model is relaxed so that it reaches hydrostatic equilibrium. In each simulation, the binary is placed on a Newtonian parabolic orbit around a $10^{6}M_{\odot}$ SMBH from a distance of $(2-3)\,r_{\rm tide}^{b}$. Note that the probability of stellar collision depends on the orbital phase of the binary. To ensure collision, we adjust the initial binary orbital phase by evolving the binary for certain duration before releasing it on the orbit toward the SMBH.
For the penetrating factor $\beta_{b}$, we select $\beta_{b}=0.6$ to represent a gentle encounter ($\beta_{b}<1$) and $\beta_{b}=5$ for a deep encounter ($\beta_{b}>1$). Throughout the paper, we consider initially coplanar, circular binaries with $\alpha=8$ (see Paper I for the effect of binary eccentricity on the collision probability).
@@ 2.2 Identify Post-collision Outcomes
In Paper I, we adopted the simplest assumption for the collision process, treating it as a perfectly inelastic collision when the separation between the two stars becomes smaller than $R_{\rm col}$, resulting in a single collision remnant. However, in our SPH simulations, we observe that collisions can lead to some degree of mass loss, and it is also possible that the two stellar remnants do not merge. From the wide range of simulations we have carried out, we select four that are representative for detailed presentation in Section 3 (see Table 1).
@@
    Head-on collision in a deep encounter ($\beta_{b}=5$). The stars collide then the binary reaches the pericenter to the SMBH. If the collision occurs with low relative velocity and a small impact parameter, the stars may merge, forming a new, more massive star.
@@
    Grazing collision in a deep encounter ($\beta_{b}=5$). In a deep encounter, the binary is disrupted after the pericenter passage. If the stars collide with a large impact parameter without merging, no further collisions occur. In this case, one star remains bound to the SMBH, while the other is ejected, similar to the Hills mechanism. However, unlike a typical binary breakup, both stars may experience violent disturbances during the process.
@@
A qualitatively description of these simulation outcomes (e.g., whether a collision occurs or how many remnants exist) is relatively straightforward. However, when performing a quantitative analysis (e.g., to determine the mass loss), careful consideration is required to distinguish SPH particles belonging to the remnant and those contributing to mass loss, particularly due to the exchange of particles between the stars during the collision.
In the case where there is only a single remnant, this analysis is relatively straightforward. We can identify a set of the SPH particles, denoted as $S_{r}$, that can be treated as a coherent body (‘remnant’). The mass, position and velocity of the remnant are
${\boldsymbol{x}}_{\rm CM}=\frac{\sum m_{i}{\boldsymbol{x}_{i}}}{\sum m_{i}},$ (3)
${\boldsymbol{v}}_{\rm CM}=\frac{\sum m_{i}{\boldsymbol{v}_{i}}}{\sum m_{i}},$
$m_{\rm rem}^{(r)}=\sum_{i\in S_{r}}m_{i},$
where $m_{i}$, ${\boldsymbol{x}_{i}}$ and ${\boldsymbol{v}_{i}}$ are the mass, position, and velocity of SPH particle $i$. In our simulations, all particles have equal mass. To simplify the calculation, we sum over all particles when computing the center of mass position and velocity. This is justified because in all our simulations, the mass loss is relatively small ($\lesssim 5\%$). Moreover, the mass lost particles are nearly symmetric relative to the remnant’s center of mass, thus having a negligible impact on the results.
@@
Our method for identifying the post-encounter remnant is based on the Bernoulli constant. For each SPH particle $i$ belonging to the remnant, a specific enthalpy is calculated according to:
$h_{i}^{(r)}=\frac{1}{2}({\boldsymbol{v}}_{i}-{\boldsymbol{v}}_{\rm CM})^{2}+u_{i}+\frac{P_{i}}{\rho_{i}}+\phi_{i}^{(r)},$ (4)
@@
where $u_{i}$ and $\rho_{i}$ are its specific internal energy and density, $P_{i}$ is the pressure, and $\phi_{i}^{(r)}$ is the gravitational potential excluding the influence of SMBH. Along a streamline with no dissipation, $h_{i}^{(r)}$ should be a constant according to the Bernoulli theorem. Particles with $h_{i}^{(r)}>0$ do not belong to $S_{r}$ they are considered as the mass loss, unbound from the remnant.
If there are two merger remnants, we can use similar algorithm based on the Bernoulli constant (see Li et al., 2021). But here, for simplicity, we adopt a more straightforward approach. First, we determine the center of mass of all SPH particles, which corresponds to the binary’s center of mass. Then, within a certain region around this center (determined by the approximate distance between the two remnants from our analysis), we compute a 2D grid-based column density, ensuring the resolution is finer than $0.1R_{\star}$. Since the orbits in our simulations are all coplanar, we assume that the $z$-coordinate of the center of mass is always zero-a simplification that has been verified to be valid. We then identify the 2D coordinates of the two density peaks as the positions of the two remnants’ centers of mass. This method is efficient; however, its drawback is that we cannot precisely determine the mass loss fraction. Nevertheless, by examining the density profile around the center of mass, we can still obtain an estimate of the mass loss.
@@ 2.3 Calculation of Mass Fallback Rate
An approximate method of calculating $\dot{M}$ is as follows. Our hydrodynamical simulations allow us to calculate the binding energy $\varepsilon$ of the mass-loss particles relative to the SMBH and to determine their energy distribution $dM/d\varepsilon$. This distribution can be used to determine the mass fallback rate $\dot{M}$ as a function of time through Kepler’s third law:
$\dot{M}(t)=\frac{dM}{d\varepsilon}\frac{d\varepsilon}{dt}=\frac{2\pi}{3}(GM_{\rm BH})^{2/3}\frac{dM}{d\varepsilon}t^{-5/3}.$ (5)
@@
The orbital eccentricity, semi-major axis and semi-latus rectum are then given by
$e=\sqrt{1+\frac{2\varepsilon h^{2}}{(GM_{\rm BH})^{2}}},\quad a=-\frac{GM_{\rm BH}}{2\varepsilon},\quad p=\frac{h^{2}}{GM_{\rm BH}}.$ (7)
The pericenter distance is then
$r_{p}=\frac{p}{1+e}.$ (8)
For bound orbits ($e<1$), the eccentric anomaly ${\cal{E}}$ and the mean anomaly ${\cal{M}}$ are
@@
For hyperbolic orbits ($e>1$), the hyperbolic anomaly ${\cal{H}}$ and the mean anomaly ${\cal{M}}$ are
@@
3 Results
Runs $\beta_{b}$ $N$ $r_{12}(R_{\odot})$ $|v_{12}|(\sqrt{GM_{\odot}/R_{\odot}})$ ${b_{12}/r_{12}}$ Number of remnant(s) ${\Delta M/M_{\rm tot}}$
1a 5 $5\times 10^{5}$ $3.749$ $1.155$ $0.0104$ $1$ $0.0495$
1b 5 $1\times 10^{5}$ $3.810$ $1.171$ $0.0625$ $1$ $0.0538$
1c 5 $1\times 10^{6}$ $3.723$ $1.147$ $0.0484$ $1$ $0.0492$
2 5 $5\times 10^{5}$ $3.612$ $1.354$ $0.4885$ $2$ $\sim 0.04$
3 0.6 $5\times 10^{5}$ $3.329$ $1.046$ $0.0285$ $1$ $0.0476$
4 0.6 $5\times 10^{5}$ $2.388$ $1.200$ $0.9085$ $1$ $0.0071$
Table 1: Summary of the SPH simulations. Here $\beta_{b}=r_{\rm tide}^{b}/r_{p}$ (with $r_{\rm tide}^{b}$ the binary tidal radius given by Eq. (1) and $r_{p}$ the pericenter distance), and $N$ refers to the number of SPH particles used to represent each star, $r_{12}$ and $v_{12}$ are the relative separation and velocity just before collision, and $b_{12}\equiv|\vec{v}_{12}\times\vec{r}_{12}|/|\vec{r}_{12}|$ is the impact parameter. We also list the number of remnant(s) and the mass loss fraction $\Delta M/M_{\rm tot}$ (where $M_{\rm tot}=2m_{\star}$) at the end of the simulation. The mass loss fractions are all calculated using the Bernoulli method (see Section 2.2), except for Run 2, where it is estimated from the upper row of Figure 9.
Our representative simulations are summarized in Table 1, corresponding to the four cases discussed in Section 2.2. For Run 1, we record the results with different resolutions (i.e., the number of SPH particles) to test the consistency of the results. We also record the relative distance $|\vec{r}_{12}|$ and velocity $|\vec{v}_{12}|$ between the two stars approximately $2t_{\rm dyn}$ (where $t_{\rm dyn}\equiv\sqrt{R_{\star}^{3}/Gm_{\star}}$) before the collision and compute the impact parameter, defined as $b_{12}\equiv|\vec{v}_{12}\times\vec{r}_{12}|/|\vec{r}_{12}|$, to get a rough idea of whether the collision is head-on ($b_{12}/r_{12}\ll 1$) or grazing ($b_{12}/r_{12}\sim 1$). Additionally, at the end of the simulation, we record the number of merger remnants and the mass loss fraction computed using the method described in Section 2.2. Note that when two merger remnants are present, we use the mass radial profile centered on each remnant’s centroid to estimate the amount of mass loss.
Figure 1: The density distribution plots of Run 1a at different times for the $z=0$ slice (the orbital plane). The density values in the plot are in visualization code units. Note that the maximum value of the color bar does not represent the actual maximum value in the plot, but is instead chosen for better visualization of the debris stream’s overall structure. In the bottom row, the small inset square shows a zoomed-in version of the region inside the dashed box from the larger plot, with the color bar adjusted for better display of the central merger remnant. The white “$\times$” marks the position of the SMBH. Here, $t=0$ approximately corresponds to the moment of collision.

### v2

@@ Abstract
When binaries are injected into low-angular-momentum orbits around a central supermassive black hole (SMBH), various outcomes can occur, including binary tidal breakup, double stellar disruptions and stellar collision. We use hydrodynamical simulations to study stellar collisions triggered by binary-SMBH encounters, examining both head-on and grazing collisions in deep ($\beta_{b}=5$) and gentle ($\beta_{b}=0.6$) encounters, where $\beta_{b}$ is the ratio of the binary tidal disruption radius to the binary pericenter distance to the SMBH. Head-on collisions consistently result in appreciable mass loss ($\sim 5\%$) and a single merger remnant. Grazing collisions have varied outcomes. In gentle encounters, multiple collisions typically form a single remnant with minimal mass loss ($\lesssim 1\%$). For deep encounters, the result depends on the specific collision parameters and stellar structure: $\gamma=5/3$ polytropic stars in our simulation produced two disturbed remnants, while solar-type stars (modeled with MESA) in our deep-grazing run formed a single merger remnant in a low-velocity collision. All merger remnants feature extended envelopes, making them susceptible to partial tidal disruptions when they return to the SMBH. The morphology and orbital energy distribution of collision-induced debris differ significantly from those of tidal disruption event (TDE) debris of single stars. Approximately half of the collision-generated debris falls back onto the SMBH, exhibiting a distinct time evolution of the fallback rate. We suggest that such mass loss and fallback can generate electromagnetic flares that mimic weak TDEs.
@@
software: Phantom (Price et al., 2018), MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023), Sarracen (Harris & Tricco, 2023), Matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011).
@@ 1 Introduction
In Yu & Lai (2024) (hereafter Paper I), we used three-body scattering experiments to characterize the key dynamics of binary tidal breakups, double stellar disruptions, and stellar collisions induced by binary-SMBH encounters. We analyzed the collision probabilities of binaries with different orbital properties and concluded that the collision probability can reach a few to 10’s percent as the binary’s pericenter distance (relative to the SMBH) varies from outside the tidal sphere to within it. Additionally, we found that the impact velocities of the colliding stars are comparable to the escape velocity of the star.
@@
Our paper is organized as follows. In Section 2, we describe the setup of our numerical simulations, including the initial conditions and methodology. In Sections 3 and 4, we present the results from our representative simulations for a $\gamma=5/3$ polytrope (representing low-mass main-sequence stars with $M\lesssim 0.4M_{\odot}$) and for a solar-like main-sequence (MS) star respectively. Finally, in Section 5, we summarize our findings and comment on the results of Paper I based on our hydrodynamical simulations.
@@ 2 Method and Setup
    $\alpha\equiv a_{b}/R_{\rm col}$, where $R_{\rm col}$ is the separation between the two stars at the moment of collision. For the equal-mass binaries ($m_{1}=m_{2}$) considered in this paper, $R_{\rm col}=2R_{\star}$.
@@ 2.1 Simulation Setup
Throughout this paper, we consider circular, equal-mass binary. In the first set of simulations (Runs 1–4 in Table 1), each star is modeled as a $\gamma=5/3$ polytrope with a mass $m_{\star}=1\,M_{\odot}$ and a radius $R_{\star}=1\,R_{\odot}$. In the second set (Runs 5 and 6), we adopt more realistic stellar models corresponding to Middle-Age Main-Sequence (MAMS) stars with a mass of $1\,M_{\odot}$ and an age of $\approx 5\rm Gyr$ and a radius of $R_{\star}\approx 1.016R_{\odot}$. The model was evolved with an initial solar metallicity of $Z=0.02$ using the one-dimensional implicit stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA), version r24.08.1 (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023). For reproducibility, we also report the central properties of the star: the central density is $\rho_{c}\approx 162.8\rm{g\,cm^{-3}}$, and the central mass fractions of hydrogen and helium are $X_{C}\approx 0.309$ and $Y_{C}\approx 0.670$, respectively. The binaries in our simulations are then created by shifting the centers of two stellar models. For the subsequent hydrodynamical evolution, we adopt the gamma equation of state (EOS) with $P=(\Gamma-1)\rho u$ , where $P$, $\rho$, and $u$ are the pressure, density, and specific internal energy, respectively, and we set the index $\Gamma=5/3$. This EOS is well-suited for capturing the essential dynamics of the collision, including shock heating. Each model is relaxed in isolation so that it reaches hydrostatic equilibrium (We fix the radial entropy profile of the star during the relaxation. We stop the relaxation procedure when the ratio of kinetic to potential energy drops below $10^{-7}$ and the error in density is less than 1%). In each simulation, the binary is placed on a Newtonian parabolic orbit around a $10^{6}M_{\odot}$ SMBH from a distance of $(2-3)\,r_{\rm tide}^{b}$. Note that the probability of stellar collision depends on the orbital phase of the binary. To ensure collision, we adjust the initial binary orbital phase by evolving the binary for certain duration before releasing it on the orbit toward the SMBH.
For the penetrating factor $\beta_{b}$, we select $\beta_{b}=0.6$ to represent a gentle encounter ($\beta_{b}<1$) and $\beta_{b}=5$ for a deep encounter ($\beta_{b}>1$). Throughout the paper, we consider initially coplanar, circular binaries with $\alpha=8$ , which corresponds to $a_{b}=16R_{\odot}$. (see Paper I for the effect of binary eccentricity on the collision probability).
@@
Runs $\beta_{b}$ NN $r_{12}(R_{\odot})$ $|v_{12}|(\sqrt{GM_{\odot}/R_{\odot}})$ $b_{12}/r_{12}$} Number of remnant(s) $\Delta M/M_{\rm tot}$}
1a 5 $5\times 10^{5}$ $3.749$ $1.155$ $0.0104$ 11 $0.0495$
1b 5 $1\times 10^{5}$ $3.810$ $1.171$ $0.0625$ 11 $0.0538$
1c 5 $1\times 10^{6}$ $3.723$ $1.147$ $0.0484$ 11 $0.0492$
2 5 $5\times 10^{5}$ $3.612$ $1.354$ $0.4885$ 22 $0.0375$
3 0.6 $5\times 10^{5}$ $3.329$ $1.046$ $0.0285$ 11 $0.0476$
4 0.6 $5\times 10^{5}$ $2.388$ $1.200$ $0.9085$ 11 $0.0068$
5 5 $5\times 10^{5}$ $4.074$ $1.186$ $0.1469$ 11 $0.0720$
6 5 $5\times 10^{5}$ $3.855$ $1.012$ $0.525$ 11 $0.0159$
Table 1: Summary of the SPH simulations. Here $\beta_{b}=r_{\rm tide}^{b}/r_{p}$ (with $r_{\rm tide}^{b}$ the binary tidal radius given by Eq. (1) and $r_{p}$ the pericenter distance), and NN refers to the number of SPH particles used to represent each star, $r_{12}$ and $v_{12}$ are the relative separation and velocity just before collision, and $b_{12}\equiv|\vec{v}_{12}\times\vec{r}_{12}|/|\vec{v}_{12}|$ is the impact parameter. We also list the number of remnant(s) and the mass loss fraction $\Delta M/M_{\rm tot}$ (where $M_{\rm tot}=2m_{\star}$) at the end of the simulation. The mass loss fractions are all calculated using the iterative Bernoulli method (see Section 2.2 below). In the first set of simulations (Runs 1–4), each star is modeled as a $\gamma=5/3$ polytrope. In the second set (Runs 5, 6), each star is modeled as a MAMS star computed from MESA. For the deep, grazing cases, the contact conditions are not identical across models (see Runs 2,6). These differences plausibly contribute to the distinct outcomes alongside stellar structure.
Our representative simulations are summarized in Table 1.
@@ 2.2 Identify Post-collision Outcomes
In Paper I, we adopted the simplest assumption for the collision process, treating it as a perfectly inelastic collision when the separation between the two stars becomes smaller than $R_{\rm col}$, resulting in a single collision remnant. However, in our SPH simulations, we observe that collisions can lead to some degree of mass loss, and it is also possible that the two stellar remnants do not merge. From the wide range of simulations we have carried out, we select four for detailed presentation in Section 3 (see Table 1). These represent four general scenarios that characterize the diversity of outcomes in our parameter space.
@@
    Head-on collision in a deep encounter ($\beta_{b}=5$). The stars collide when the binary reaches the pericenter to the SMBH. If the collision occurs with low relative velocity and a small impact parameter, the stars may merge, forming a new, more massive star.
@@
    Grazing collision in a deep encounter ($\beta_{b}=5$). In a deep encounter, the binary is disrupted after the pericenter passage. High-velocity, large-impact-parameter encounters that avoid stellar merger typically lead to an outcome that is similar to the Hills mechanism of binary breakup: one star is ejected, while the other remains bound to the SMBH. A key difference is that both interacting stars may suffer significant structural disturbances during the grazing collision. In contrast, low-velocity encounters can result in a merger even at large impact parameters. In this work, whether a merger occurs depends on the specific values of $v_{12}$ and $b_{12}/r_{12}$ at contact (see Table 1), rather than the stellar structure alone. Such a merger forms a new, more massive star.
@@
A qualitative description of these simulation outcomes (e.g., whether a collision occurs or how many remnants exist) is relatively straightforward. However, when performing a quantitative analysis (e.g., to determine the mass loss), careful consideration is required to distinguish SPH particles belonging to the remnant and those contributing to mass loss, particularly due to the exchange of particles between the stars during the collision.
We employ an iterative scheme to quantitatively classify SPH particles as belonging to one or two gravitationally bound stellar cores ($S_{k}$) or as part of the unbound mass loss. The algorithm begins with an initial assignment of particles to potential cores and then iteratively refines this classification until convergence is achieved. The following describes each step of this procedure in detail.
To initiate the iterative process at a given snapshot, we first require an initial assignment of particles to potential cores. The procedure differs for single-core and dual-core cases:
  • Single-core case:
    1. 1.
      Compute the global center-of-mass, $\vec{r}_{\rm CM,global}$, using all SPH particles.
    2. 2.
      Assign all particles $i$ with $|\vec{r}_{i}-\vec{r}_{\rm CM,global}|\leq R_{\rm guess}$ as the initial members of the single core $S_{1}$.
  • Dual-core case:
    1. 1.
      Construct a 2D grid of the projected column density (since the orbits are coplanar, the center-of-mass lies in the $z=0$ plane). Ensure the grid resolution is finer than $0.1\,R_{\star}$.
    2. 2.
      Identify the two density peaks from the 2D grid as the initial core centers, denoted by $\vec{r}_{\rm CM,init}^{(1)}$ and $\vec{r}_{\rm CM,init}^{(2)}$.
    3. 3.
      For each SPH particle $i$, calculate its distances to the two centers: $d_{i}^{(1)}=|\vec{r}_{i}-\vec{r}_{\rm CM,init}^{(1)}|$ and $d_{i}^{(2)}=|\vec{r}_{i}-\vec{r}_{\rm CM,init}^{(2)}|$.
    4. 4.
      Assign particle $i$ to core $S_{1}$ if $d_{i}^{(1)}\leq R_{\rm guess}$ and either $d_{i}^{(2)}>R_{\rm guess}$ or $d_{i}^{(1)}\leq d_{i}^{(2)}$.
    5. 5.
      Otherwise, assign particle $i$ to core $S_{2}$ if $d_{i}^{(2)}\leq R_{\rm guess}$ and either $d_{i}^{(1)}>R_{\rm guess}$ or $d_{i}^{(2)}<d_{i}^{(1)}$.
The parameter $R_{\rm guess}$ can be adjusted to optimize the initial identification performance; we adopt $R_{\rm guess}=1R_{\star}$ for single-core cases and $R_{\rm guess}=2R_{\star}$ for dual-core cases.
Once the cores are initialized, the main iterative loop begins. In each iteration, we first update the core properties. For every identified core $k$ (with $k=1$ for the single-core case, and $k=1,2$ for the dual-core case), we first compute the following quantities based on the current set of constituent particles $S_{k}$:
  • Number of particles ($N_{k}$): The number of SPH particles currently assigned to core $S_{k}$. Given equal particle mass $m_{p}$, the total mass of the remnant is $M_{k}=N_{k}\times m_{p}$.
  • Center-of-mass position $\vec{r}_{\rm CM}^{(k)}$ and velocity $\vec{v}_{\rm CM}^{(k)}$ of the core:
    $\vec{r}_{\rm CM}^{(k)}=\frac{1}{N_{k}}\sum_{i\in S_{k}}\vec{r}_{i},\quad\vec{v}_{\rm CM}^{(k)}=\frac{1}{N_{k}}\sum_{i\in S_{k}}\vec{v}_{i}.$ (3)
@@
Our method for identifying the post-encounter remnant is based on the Bernoulli constant. For each SPH particle $i$ belonging to the remnant $k$, a specific enthalpy is calculated according to:
$h_{i}^{(k)}=\frac{1}{2}(\vec{v}_{i}-\vec{v}^{(k)}_{\rm CM})^{2}+u_{i}+\frac{P_{i}}{\rho_{i}}+\phi_{i}^{(k)},$ (4)
@@
where $u_{i}$ and $\rho_{i}$ are its specific internal energy and density, $P_{i}$ is the pressure, and $\phi_{i}^{(k)}$ is the gravitational potential due to the mass distribution of core $k$ excluding the influence of SMBH. Along a streamline with no dissipation, $h_{i}^{(k)}$ should be a constant according to the Bernoulli theorem. Particles with $h_{i}^{(k)}>0$ do not belong to $S_{k}$ - they are considered as the mass loss, unbound from the remnant.
The kinetic and internal energy terms in Equation (4) are readily available from the simulation output. However, the evaluation of $\phi_{i}^{(k)}$ is more challenging. While the most accurate approach would be a direct summation of the potential contributions from all particles $j\in S_{k}$, this method is computationally prohibitive for large $N$. Therefore, we adopt two approximate methods for estimating $\phi_{i}^{(k)}$, tailored for the single-core and dual-core cases:
  • Global-approximation for the single-core case:
    This approach estimates the gravitational potential from the remnant by subtracting the known SMBH contribution $\phi_{\text{BH},i}$ from the total potential at particle $i$:
    $\phi_{i}^{(k)}\simeq\phi_{\text{total},i}-\phi_{\text{BH},i},$ (5)
    where $\phi_{{\rm total},i}$ is provided by the SPH simulation output.
  • Point-mass-approximation for the dual-core case:
    Here, core $k$ is modeled as a point mass $M_{k}=N_{k}m_{p}$ located at its center-of-mass $\vec{r}_{\rm CM}^{(k)}$. Thus,
    $\phi_{i}^{(k)}\approx-G\frac{M_{k}}{|\vec{r}_{i}-\vec{r}_{CM}^{(k)}|}.$ (6)
    To avoid singularities, a softening length or a lower bound on $|\vec{r}_{i}-\vec{r}_{\rm CM}^{(k)}|$ (e.g., $\epsilon$) is applied.
After calculating the specific enthalpy $h_{i}^{(k)}$, we re-assign each particle according to the following criteria:
  • Single-core case: If $h_{i}^{(1)}<0$, particle $i$ is considered gravitationally bound and is assigned to (or retained in) core $S_{1}$. Otherwise, it is classified as unbound.
  • Dual-core case: Particle $i$ is assigned to core $S_{k}$ (with $k=1$ or $2$) for which $h_{i}^{(k)}$ is minimized, provided that the minimum $h_{i}^{(k)}<0$. If $h_{i}^{(k)}\geq 0$ for all cores, the particle is considered unbound.
The sets of particles $\{S_{k}\}$ are then updated based on the new assignments, and the process is repeated until convergence is achieved.
Although most of our calculations converge within a few iterations, we implement an additional optimization to further accelerate convergence during the later stages, when the cores are already well established and changes in membership occur primarily at the outskirts of the remnant. After a specified number of initial full iterations, particles currently assigned to core $k$ and located within a small radius $R_{\rm frozen}$ (e.g., $R_{\rm frozen}=R_{\star}$) of that core’s center-of-mass $\vec{r}_{\rm CM}^{(k)}$ are designated as frozen. These particles are assumed to remain bound and are excluded from subsequent enthalpy evaluations and re-assignment steps. In each following iteration, only particles outside the frozen region or those previously classified as unbound are re-evaluated. Nevertheless, the center-of-mass position and total mass (or particle count) of each core are still computed using the complete set of its current members, including both frozen and re-assigned particles. This strategy significantly reduces the number of particles requiring costly enthalpy calculations at each step, thereby improving computational efficiency without sacrificing accuracy.
Once the iterative process converges, any particle not assigned to a bound core ($S_{k}$) is classified as mass loss. For subsequent analysis, we record the indices of all SPH particles, distinguishing between those assigned to bound cores and those identified as unbound.
@@ 2.3 Calculation of Mass Fallback Rate
An approximate method of calculating $\dot{M}$ is as follows. Our hydrodynamical simulations allow us to calculate the specific binding energy $\varepsilon$ of the mass-loss particles relative to the SMBH and to determine their energy distribution $dM/d\varepsilon$. This distribution can be used to determine the mass fallback rate $\dot{M}$ as a function of time through Kepler’s third law:
$\dot{M}(t)=\frac{dM}{d\varepsilon}\frac{d\varepsilon}{dt}=\frac{(2\pi GM_{\rm BH})^{2/3}}{3}\frac{dM}{d\varepsilon}t^{-5/3}.$ (7)
@@
The semi-major axis, orbital eccentricity and pericenter distance are then given by
$\quad a=-\frac{GM_{\rm BH}}{2\varepsilon},e=\sqrt{1-\frac{h^{2}}{GM_{\rm BH}a}},\quad r_{p}=\frac{h^{2}}{GM_{\rm BH}(1+e)}.$ (9)
To determine the angle $\theta$ between the particle’s instantaneous position and its pericenter direction, we compute the eccentricity vector
$\mathbf{e}=\frac{\mathbf{v}\times\mathbf{h}}{GM_{\rm BH}}-\frac{\mathbf{r}}{r},$ (10)
which points from the SMBH toward the orbital pericenter. The angle $\theta$ is then obtained via
$\cos\theta=\frac{\mathbf{e}\cdot\mathbf{r}}{er}.$ (11)
[…]
@@
[…]
@@
3 Results: $\gamma=5/3$ polytrope
Here we present our numerical results for $\gamma=5/3$ polytrope stars, representing $M\lesssim 0.4M_{\odot}$ MS stars.
For Run 1, we record the results with different resolutions (i.e., the number of SPH particles) to test the consistency of the results. We also record the relative distance $|\vec{r}_{12}|$ and velocity $|\vec{v}_{12}|$ between the two stars approximately $2t_{\rm dyn}$ (where $t_{\rm dyn}\equiv\sqrt{R_{\star}^{3}/Gm_{\star}}$) before the collision and compute the impact parameter, defined as $b_{12}\equiv|\vec{v}_{12}\times\vec{r}_{12}|/|\vec{v}_{12}|$, to characterize the nature of the encounter. Throughout this paper, we define a collision as “nearly head-on” if the impact parameter is a small fraction of the contact separation ($b_{12}/r_{12}\lesssim 0.15$), and as “grazing” if the impact parameter is a significant fraction of it ($b_{12}/r_{12}\gtrsim 0.5$). For brevity, we will hereafter refer to “nearly head-on” collisions simply as “head-on” collisions. Additionally, at the end of the simulation, we record the number of merger remnants and the mass loss fraction computed using the method described in Section 2.2. Note that when two merger remnants are present, we use the mass radial profile centered on each remnant’s centroid to estimate the amount of mass loss.
Figure 1: The density distribution plots of Run 1a at different times for the $z=0$ slice (the orbital plane). The density values in the plot are in visualization code units. Note that the maximum value of the color bar does not represent the actual maximum value in the plot, but is instead chosen for better visualization of the debris stream’s overall structure. In the bottom row, the small inset square shows a zoomed-in version of the region inside the dashed box from the larger plot, with the color bar adjusted for better display of the central merger remnant. The white $\times$” marks the position of the SMBH. Here, $t=0$ approximately corresponds to the moment of collision.